OpenStudy (akashdeepdeb):

Help needed. Find the vector sum of the forces \(\vec{A}\) and \(\vec{B}\). \(\vec{| A |}\) is 130N and is along \(\vec{OA}\) where A is (3,-4,12) and \(\vec{| B |}\) = 200N and is along vector \(\vec{PQ}\) where \(\vec{P}\) is (-2,4,7) and Q is (-2,8,10). DO NOT SOLVE IT. Only tell me the concept here. :D

OpenStudy (anonymous):

Draw a coordinate system and then put vectors on ...

OpenStudy (akashdeepdeb):

Yes I tried. Can you show it to me once?

OpenStudy (akashdeepdeb):

@Saeeddiscover ?

OpenStudy (akashdeepdeb):

@.Sam. ? @experimentX ? :D

OpenStudy (anonymous):

To add vectors, you need to put them at a common point without changing their directions.The vector OA makes no difficulty, you may calculate the direction of vector PQ in space, and then catch it to put in the origin.

OpenStudy (akashdeepdeb):

Why have they specified their Magnitudes? And what is O in OA?

OpenStudy (anonymous):

O just refers to the point (0,0,0). you might calculate the directional angles to find its direction into space...

OpenStudy (akashdeepdeb):

And why are the magnitudes given? I mean How is the magnitude of A = 120N ?

OpenStudy (anonymous):

Each of the coordinate axes are written in terms of newton rather than meter.

OpenStudy (akashdeepdeb):

No I mean Is \(\vec{A}\) = 3i - 4j + 12k ? If yes then HOW IS ITS MAGNITUDE 120?

OpenStudy (akashdeepdeb):

130 sry.

OpenStudy (akashdeepdeb):

@Saeeddiscover ?

OpenStudy (anonymous):

The magnitude of any vector can be obtained using the relation |A|=(a1^2+a2^2+...+an^2)1/2 But in this example, we obtain sth different. Maybe you need to scale the coordinates ...

OpenStudy (akashdeepdeb):

Yes man I know that! It is just that after I calculated the displacement vector for PQ I am not getting the magnitude as 200N And even after scaling OA I did not get 130N, What seems to be the problem here chief? :D

OpenStudy (akashdeepdeb):

?

OpenStudy (anonymous):

For OA and PQ, we have: |OA|=(3^2+4^2+12^2)^1/2=13 |PQ|=[(-2--2)^2+(8-4)^2+(10-7)^2]^1/2=5 If we multiply OA in 10 we get 130.....

OpenStudy (akashdeepdeb):

And is it the same for PQ?

OpenStudy (akashdeepdeb):

It is NOT! So now what? These has to be a reason they gave the magnitudes! :/

OpenStudy (akashdeepdeb):

@Saeeddiscover ?

OpenStudy (akashdeepdeb):

@thomaster ?

OpenStudy (akashdeepdeb):

@Yahoo! HELP? No help on YahooAnswers too! ^_^ :D Haha..My corny puns! :D

OpenStudy (akashdeepdeb):

@Preetha ? @terenzreignz ? @thomaster ?

OpenStudy (akashdeepdeb):

@radar ? @goformit100 ? @somebodyhelpmeee.... :/

OpenStudy (akashdeepdeb):

@AravindG ?

OpenStudy (anonymous):

wat u really want here ?

OpenStudy (akashdeepdeb):

The vector sum of the 2 vectors! :/

OpenStudy (akashdeepdeb):

Please tell me the question is wrong!! :/ I ahve been at it for a long time

OpenStudy (anonymous):

No,I think that the question is correct :P ... but I can't solve "3D coordinate system" and i'm not very good at it,sorry :(

OpenStudy (akashdeepdeb):

It is fine. Thanks Stephen Hawking! :D

OpenStudy (anonymous):

i dont thik so hw will |B| = 200N?

OpenStudy (akashdeepdeb):

The book says it. :/ Do you think it should be 50N?

OpenStudy (anonymous):

wat shuld be ur answer?

OpenStudy (anonymous):

i mean the net force

OpenStudy (akashdeepdeb):

sqrt 234.

OpenStudy (anonymous):

the resultant vector b/w two vectors at an angle theta is V= sqrt(A^2 + B^2 +2AB cos theta)

OpenStudy (anonymous):

i think u have to find cose theta form given expression and subs

OpenStudy (akashdeepdeb):

No actually I just add the 2 vectors! And then I got 3i + 15k

OpenStudy (akashdeepdeb):

Sorry....What I mean to say is that... I just added the two vectors with their vector components OA + PQ and then I got the vector = 3i + 15k So then I got the magnitude as sqrt 234!! :D

OpenStudy (akashdeepdeb):

@Yahoo! I don't get it. :/

OpenStudy (experimentx):

first find the vectors and just add it up

OpenStudy (akashdeepdeb):

I did. Can you PLEASE just check if my answer is right? THERE IS NO SOLUTION IN THIHS BOOK.

OpenStudy (akashdeepdeb):

The thing is how do I find OA? Do I take O as 0? Should O actually be Q? [Considering it is a typo?]

OpenStudy (experimentx):

A = 130*(3,-4,12)/sqrt(3^2 + 4^2 + 12^2) = 130*A/|A| B = 200 * (Q - P)/|Q-P| then just add them up

OpenStudy (akashdeepdeb):

The question is why are you multiplying the given magnitudes by their ACtual magnitudes?

OpenStudy (experimentx):

you are given direction ... to get vector multiply the magnitude by direction.

OpenStudy (akashdeepdeb):

But then you are dividing also right?

OpenStudy (experimentx):

what is the vector along (1,1,1) whole magnitude is 12?

OpenStudy (akashdeepdeb):

Okay so the vector would be i + j +k But direction would be 1/12.1 + 1/12.j + 1/12.k ??

OpenStudy (experimentx):

No ...

OpenStudy (akashdeepdeb):

Then?

OpenStudy (experimentx):

the first thing that comes to your mind is 12*(i + j + k) = 12 i + 12 j + 12 k but if you take its magnitude ... it turns out to be sqrt(12^2 + 12^2 + 12^2) = 12 sqrt(3)

OpenStudy (experimentx):

whne you are given direction first, you should normalize it. that means you should find unit vector along that direction.

OpenStudy (experimentx):

(1,1,1) is not unit vector. to find unit vector along (1,1,1) you need to divide it by it's magnitude. (1,1,1)/|(1,1,1)| = 1/sqrt(3)(1,1,1) <- now this is unit vector.

OpenStudy (experimentx):

then multiply it by 12, you get a vector whose magnitude is 12 and direction is along (1,1,1)

OpenStudy (akashdeepdeb):

Multiply this by 12 [1/sqrt(3)(1,1,1)]

OpenStudy (akashdeepdeb):

?

OpenStudy (experimentx):

yes

OpenStudy (akashdeepdeb):

What is (1,1,1) then?

OpenStudy (experimentx):

this is just given direction.

OpenStudy (akashdeepdeb):

But it is just a POINT.

OpenStudy (experimentx):

Oh ... you are not familiar with tuple notation of vector. (1,1,1) = i + j + k

OpenStudy (akashdeepdeb):

Not quite.... We have not studied 3D Geometry yet. :/

OpenStudy (experimentx):

anyway you got the point right.

OpenStudy (akashdeepdeb):

Yes pretty much! And If we are given just one point does that mean that It starts from the origin?

OpenStudy (experimentx):

no not quite.

OpenStudy (experimentx):

it has to be specified that other end is at origin.

OpenStudy (experimentx):

as in your question. I didn't do it becuase i am lazy to type.

OpenStudy (akashdeepdeb):

I got that... :D Thanks btw for helping. Should I be knowing 3D Geometry to solve this question?

OpenStudy (experimentx):

no .. just use your i,j,k notation instead of (a,b,c)

OpenStudy (akashdeepdeb):

Yeah okay. Thanks! :)

OpenStudy (experimentx):

fist make the vector OA = 3i-4j+12k then make it unit vector (unit OA)=OA/|OA| then multiply it by magnitude = 130 *(unit OA) then you are done, you can do what is specified.

OpenStudy (akashdeepdeb):

Okay.

OpenStudy (anonymous):

Have you been satisfied?

OpenStudy (akashdeepdeb):

98%

OpenStudy (anonymous):

What's still been unsolved?

OpenStudy (aravindg):

Sorry I was away at that time.

OpenStudy (akashdeepdeb):

It is totally alright! :)